Worked out problems: Solve completely algebraically before inserting any values. Make sure your algebraic answer is in terms of given variables 5. A rock is thrown over a cliff from the origin with initial velocity +z points in the upward direction. where is it a time later? Ve,+o where A. (4 points) List two assumptions that you need to make to solve this projectile motion problem. B. (8 points) For the initial velocity, = 3 Isi-24, +2.02; where is the rock 2.0 s later? C. (8 points) If the ground is 65 m below the cliff, when will the rock hit the ground? Use the same initial velocity as in part B. Need some help.
Answer #1

5) XY plane is horizontal and Z-axis is pointing upward.

Initial velocity of rock \vec{v}_0=v_{ox}\hat{x}+v_{oy}\hat{y}+v_{oz}\hat{z}

After time \Delta t ,

The X-coordinate of rock is x(t)=v_{ox}\Delta t+\frac{1}{2}a_x(\Delta t)^2

The Y-coordinate of rock is y(t)=v_{oy}\Delta t+\frac{1}{2}a_y(\Delta t)^2

The Z-coordinate of rock is z(t)=v_{oz}\Delta t+\frac{1}{2}a_z(\Delta t)^2

At time \Delta t , the position of particle is \vec{r}=\vec{v}_0 \Delta t +\frac{1}{2}\vec{a}(\Delta t)^2

A) The two assumptions are,

Acceleration is constant and acting vertically downwards , There is no air resistance in the rock's motion.

Since X and Y-axes are horizontal (and perpendicular to each other) , a_x=0 , a_y=0 and a_z=-g

So the velocity along X-axis and velocity along Y-axis do not change with time.

Velocity along Z-axis changes with time since the acceleration along Z-axis is changing.

Position coordinates of the rock after time \Delta t,

x=v_{ox}\Delta t , y=v_{oy}\Delta t ,  z=v_{oz}\Delta t-\frac{1}{2}g(\Delta t)^2

B)

\vec{v}_0=(3.1\hat{x}-2.4\hat{y}+2.0\hat{z})\,m/s\Delta t=2.0\,s

x=v_{ox}\Delta t=3.1*2.0=6.2\,m

y=v_{oy}\Delta t=-2.4*2.0=-4.8\,m

z=v_{oz}\Delta t-\frac{1}{2}g(\Delta t)^2=2.0*2.0-\frac{1}{2}*9.81*(2.0)^2=-15.62\,m

(x,y,z)=(6.2,-4.8,-15.62)\,m

\vec{r}=(6.2\hat{x}-4.8\hat{y}-15.62\hat{z})\,m

C) Given that Z=-65\,m ( downward vectors are taken negative)

z=v_{oz}\Delta t-\frac{1}{2}g(\Delta t)^2

-65=2.0\,t-\frac{1}{2}*9.81 \,t^2

4.904\,t^2-2\,t-65=0

solving above quadratic equation t=3.9\,s